OpenStudy (anonymous):

I have tried this integral but I can't get it right 2 sin y^2 dy from (pi/2)^1/2 to x. Can you help me?

OpenStudy (dayakar):

type what u did

OpenStudy (anonymous):

It's too long but I used u substitution for x^2, so I got sen u du/2y

OpenStudy (thomas5267):

Is this the integral? \[ \int_{\left(\frac{\pi}{2}\right)^{1/2}}^x 2\sin^2\left(y^2\right)\,dy \] Is this question asking for something like this? \[ \frac{d}{dx}\int_{\left(\frac{\pi}{2}\right)^{1/2}}^x 2\sin^2\left(y^2\right)\,dy \]